# CER-HV: A CER-Based Human-in-the-Loop Framework for Cleaning Datasets Applied to Arabic-Script HTR

### v1

@@ Abstract
handwritten text recognition, data cleaning, Arabic scripts, label noise detection, CRNN, Pashto, Urdu, Persian, Ajami
@@
[label1] organization=Luleå University of Technology, Department of Computer Science, Electrical and Space Engineering, Embedded Intelligent Systems LAB, city=Luleå, postcode=97187, country=Sweden
@@ 1 Introduction
Handwritten text documents are a key part of cultural heritage and preserve historical knowledge across generations. Handwritten Text Recognition (HTR) aims to automatically convert images of handwritten documents into machine-readable text. This supports the preservation, search, and analysis of valuable materials and enables applications such as digitizing of archives or assisting individuals with visual impairments [11].
This work focuses on the recognition of handwritten text at the line level, where the goal is to predict the character sequence contained in a single extracted line of text. Recent work highlights this level of annotation as a practical and widely adopted choice in HTR, offering a balance between annotation effort and the ability to capture meaningful contextual information [18]. Compared to character- or word-level tasks, line-level recognition is more challenging and better reflects real document conditions [11].
HTR for Latin-script languages has reached near-human performance, especially for English [11]. In contrast, the same level of success has not been achieved for Arabic-script languages, which are used by more than 400 million people worldwide [30]. This gap is often attributed to the characteristics of Arabic script, which is cursive in nature, many letters differ only by small diacritical marks, and their shapes change depending on their position within a word, and available datasets are fewer, smaller, and less diverse than Latin-script resources [31]. To reduce this performance gap, research has mostly focused on architectural improvements while paying much less attention to data quality [6, 2].
@@
These issues do not arise randomly; they often originate from the semi-automated way in which handwritten text datasets are constructed, where automated methods are combined with human oversight. Because producing large datasets is time and labor intensive, this oversight is not always adequately managed, allowing samples with errors—such as incorrect or incomplete transcriptions, segmentation problems (e.g., truncated lines or multiple text lines within a single image), orientation mistakes, script-mismatched samples, or non-textual document elements (e.g., stamps or signatures)—to be included in the final dataset. Label errors are problematic as deep neural networks can memorize noisy labels during training [41], which harms generalization [35] and can distort model selection and benchmarking [24].
@@
However, automated scoring alone can mistakenly flag valid but difficult samples as noisy, as prior work has shown that only about half of algorithmically flagged candidates correspond to genuine label errors [24]. Given the relatively small size of typical HTR datasets (e.g., IAM $\approx 10K$ lines [22], KHATT $\approx 6K$ [20]), human inspection of high-CER samples is practical. We therefore introduce CER-based Ranking with Human Verification (CER-HV), as a two-stage framework that combines CER-based ranking with human verification. To our knowledge, no prior work has combined noise detection with human verification in sequence-to-sequence (Seq2Seq) settings such as HTR.
@@
    We establish a strong CRNN baseline for Arabic-script HTR that achieves state-of-the-art results on multiple datasets, including a reduction from 20.7% to 8.26% CER on Pashto (PHTI), and the first benchmarks for Ajami and Persian line-level HTR.
@@
    We provide the first publicly available cleaned evaluation splits and line-level benchmarks for Persian handwritten text (PHTD), including extracted line images and test partitions, enabling reproducible comparison and future benchmarking.
@@ 2.1 Arabic-script HTR
A similar direction is observed in Urdu HTR, where Transformer-based architectures have also been explored [28], often relying on additional printed or synthetic text data to supplement handwritten training sets due to their data-intensive nature [21, 14, 13].
@@ 2.2 Label Error Detection in Handwritten Text Recognition
Another family of methods exploits learning dynamics to identify noisy labels. O2U-Net ranks samples based on their average training loss over cycles of underfitting and overfitting, under the observation that mislabeled samples tend to maintain higher loss throughout training [15]. This idea is closely related to the memorization effect, where deep networks fit clean samples earlier than noisy ones. This early-learning phenomenon has been analyzed more broadly in the context of learning with noisy labels, where several works show that deep networks first fit clean labels before memorizing noisy ones, and propose training strategies that explicitly exploit or regularize this behavior to improve robustness [19, 38]. While O2U-Net is formulated for classification, its reliance on per-sample loss trajectories makes it conceptually applicable to other settings. However, in HTR, the interpretation of per-sample CTC loss is complicated by alignment variability and structural errors, and cyclic overfitting strategies may not always be desirable. Prior work on early stopping shows that preventing late-stage memorization can be beneficial under label noise, suggesting that careful control of training dynamics is important when adapting such methods [34].
@@ 3.1 Problem Formulation
$P(y\mid x;\theta)=\sum_{\pi\in(\Sigma^{\prime})^{T},B(\pi)=y}\prod_{t=1}^{T}P(\pi_{t}\mid x;\theta)$ (2)
and the model parameters $\theta$ are learned by minimizing the CTC loss over the training set
@@ 3.2 HTR Model: CRNN Architecture
The CRNN model performs handwritten text line recognition using CTC [12], eliminating the need for explicit character-level segmentation. Given an input image of size H×W (height × width), the network interprets it as a sequence of W narrow vertical slices, each representing a small visual context of the image. The model outputs a sequence of character probability distributions of length L, typically shorter than W due to spatial downsampling. During decoding, consecutive duplicate predictions are merged and blank tokens are removed to yield the final transcription.
@@
Figure 3: The CRNN architecture used as the base recognition model in Stage 1 of the CER-HV framework. The model combines a convolutional encoder for visual feature extraction with an RNN decoder and a CTC output layer for sequence prediction.
The CNN encoder extracts hierarchical visual features from the input text line image, producing a sequence of 2D feature maps. An intermediate column-wise max-pooling step reduces the feature map height to one, effectively converting the 2D representation into a 1D sequence of feature vectors along the image width. This operation preserves the horizontal spatial order while reducing sensitivity to absolute vertical position, preparing the sequence for the recurrent decoder.
The RNN decoder, composed of three stacked bidirectional LSTM layers (256 hidden units each), models contextual dependencies across time steps in both directions. Each recurrent output is passed through a fully connected classifier layer that maps it to a vector of dimension equal to the character set size (including the CTC blank token). The resulting per-frame character probability distributions are then decoded by the CTC layer to produce the final recognized text sequence without requiring explicit segmentation.
@@
Concretely, the CNN encoder begins with a 7×7 convolution (32 channels, stride 2×2), followed by ResNet blocks (2 with 64 channels, 3 with 128, and 2 with 256), interleaved with batch normalization and dropout. Progressive 2×2 max-pooling between block groups downsamples the spatial resolution, producing feature maps of size (H/8) × (W/8) × 256. Column-wise max pooling then flattens this tensor into a sequence of (W/8) feature vectors (256 dimensions each).
@@ 3.3 CER-HV Framework: CER Based Ranking and HITL Verification
Although automated scoring highlights suspicious samples, some correctly labeled but visually difficult examples may also receive high CER. For this reason, a HITL verification step is used, in which samples whose CER exceeds a threshold are selected for review:
@@
After human verification and cleaning, the cleaned dataset $D^{\prime}$ is used to retrain the model. The retrained recognizer is evaluated on the test set, and the same CER based scoring procedure can be used to identify remaining issues in the evaluation data. The process may be repeated if needed. Algorithm 1 summarizes the complete procedure. Beyond cleaning existing benchmarks, CER-HV can be applied as a general validation procedure to assess label quality in newly created or previously unverified line-level HTR datasets. Training details, hyperparameters, input preprocessing, and experimental settings are described in Section 4.
@@
 if label is incorrect then
    Relabel as $(x_{i},\tilde{y}_{i})$;
  else if sample has segmentation, orientation, or irrelevant content then
@@ 4 Experimental Setup
This section describes the datasets and training setup used to evaluate the CER-HV framework.
@@ 4.1 Datasets
This study employs six Arabic-script HTR datasets covering five languages segmented and labeled at the line level, covering both modern and historical handwriting. Together, these datasets provide broad coverage of writing styles, domains, and regional variations across Arabic, Persian, Urdu, Pashto, and West African Ajami scripts. Examples from the datasets are shown in Figure 4, and dataset statistics are summarized in Table 1. For datasets without public train, validation, and test splits (e.g., PHTD and Ajami), we applied an 80%, 10%, and 10% division. These datasets collectively enable assessment of the robustness of the proposed framework across diverse handwriting styles and linguistic varieties.
@@
KHATT (Arabic): The KHATT dataset [20] is a widely used benchmark for modern Arabic handwriting recognition. It contains paragraph images written by 1,000 writers from diverse demographic backgrounds. Each writer contributed both repeated and unique texts under controlled imaging conditions (black ink on white background). In our experiments, we used only the unique-text subset, yielding 6,742 segmented line images.
@@
NUST-UHWR (Urdu): The NUST-UHWR dataset [36] is the first published large-scale, unconstrained Urdu handwriting corpus. It includes 10,608 text lines written by nearly 1,000 contributors. Text samples span seven topical domains to ensure linguistic variety. The dataset is partitioned into 8,483 training, 1,061 validation, and 1,061 test lines.
@@ 4.2 Training Setup and Evaluation
For each dataset, image preprocessing followed a standardized resizing and padding procedure to reduce shape distortion and ensure uniform input dimensions. Specifically, all line images were resized to the rounded average height and width (in whole pixels) of their respective training set, as reported in Table 1. If an image exceeded these target dimensions, it was proportionally rescaled; otherwise, it was centered and padded with median pixel values. This approach preserved the aspect ratio while providing consistent spatial context across samples, which we observed stabilized training and improved recognition performance on longer text lines. In addition to resizing, we applied fixed horizontal padding of 64 pixels to the left and right sides of each line image, following a preprocessing strategy reported in prior work [7], which demonstrated improved recognition performance in CRNN-based line-level HTR.
To further regularize training and enhance robustness to handwriting variability, we applied a consistent set of data augmentations across all datasets. These include affine transformations (small rotation and skew), grid distortion, elastic deformation, Gaussian noise addition, brightness and contrast adjustment, and morphological operations (erosion and dilation). Such augmentations have been shown to improve generalization in HTR tasks by encouraging invariance to writer style, line curvature, and ink intensity.
For each dataset, a separate CRNN model was trained for a maximum of 800 epochs with a batch size of 16 using the Adam optimizer and an initial learning rate of $5\times 10^{-4}$. A multi-step learning rate scheduler reduced the learning rate by a factor of 0.1 at 50% and 75% of the total epochs. Early stopping with patience $P=20$ epochs was applied based on validation CER to prevent overfitting. For each dataset, the model achieving the lowest CER on the validation set was selected for final evaluation on the test partition.
@@ 5 Results
We begin by comparing our CRNN results with prior work across all datasets using the original, unaltered data to establish quantitative CER baselines. We then apply the CER-HV framework to analyze label noise across the training, validation, and test splits. Representative qualitative examples of common labeling issues are presented to illustrate the identified error categories. Finally, we quantify the impact of correcting these errors by evaluating how cleaning the affected splits influences recognition performance.
@@ 5.1 Quantitative Benchmarking of the CRNN-Based HTR System
Table 2: Comparison of our CRNN with established methods from the literature. Datasets, models, architectures, synthetic data usage, and CER results across datasets (reported as mean ± standard deviation). A dashed line separates non-Transformer architectures from Transformer-based models.
Dataset & Language Model Architecture Synthetic Data CER (%)
KHATT (Arabic)[20] [30] SFR Yes 14.1
[5] CNN+BiLSTM+CTC No 15.45
[5] EfficientNetB4BLSTM Yes 13.34
[39] FCN No 8.77
Ours CRNN No 8.45 ± 0.14
[23] Transformer with cross-attention Yes 18.45
[6] TrOCR Yes 15.4
[2] DAN No 8.9
Muharaf (Arabic) [30] [30] SFR Yes 18.1
Ours CRNN No 10.11 ± 0.055
[6] TrOCR Yes+11K restricted data 11.7
PHTI (Pashto) [16] [17] MD-LSTM No 20.7
Ours CRNN No 8.26 ± 0.05
PHTD (Persian) [1] Ours CRNN No 11.3 ± 0.07
NUST-UHWR (Urdu) [36] [36] Modified CRNN 7.35
[13] ET-Network(No Attention) No 7.79
Ours CRNN No 6.66 ± 0.18
[14] ViT+T5+GPT2 No 17.0
[21] CNN-Transformer (pre-trained) Yes 6.2
[13] ET-Network(with Attention) Yes 5.70
[28] Conv-Transformer No 6.4
Ajami (Fulfulde/Hausa) [40] [40] Pre-trained models from OPEN ITI No 64–84
Ours CRNN No 10.66± 0.04
@@ 5.2 Label Error Analysis
Using the CER-HV process described in Section 3.3, we analyzed label and content errors across all six datasets and across all dataset splits (training, validation, and test). The analysis reveals that the presence, type, and prevalence of labeling issues vary substantially depending on both the dataset and the split. In total, five datasets contained mislabeled or corrupted samples. PHTD did not exhibit any detectable labeling errors and is therefore excluded from the subsequent error analysis. For the remaining datasets, entries marked with “—” indicate that the corresponding error type was not observed for that dataset and spli
@@
Table 3: Precision of the CER-based noise detection computed over the top 50 highest-ranked samples in each split for datasets with sufficient numbers of flagged samples. Precision reflects the proportion of flagged samples that were confirmed as true label errors; the remaining cases correspond to valid but difficult samples.
@@
Table 4: Distribution of detected label errors by type across training, validation, and test splits based on the CER-HV annotation process.
Dataset Split
Transcription
Error
Segmentation
Error
Orientation
Error
Script-
Mismatch
Irrelevant
Content

### v2

@@ Abstract
handwritten text recognition, label noise detection, CRNN, Pashto, Urdu, Persian, Ajami
@@
[label1] organization=Luleå University of Technology, Department of Computer Science, Electrical and Space Engineering, city=Luleå, postcode=97187, country=Sweden
@@ 1 Introduction
Handwritten text documents preserve cultural heritage and historical knowledge across generations. Handwritten Text Recognition (HTR) aims to automatically convert images of handwritten documents into machine-readable text. This supports the preservation, search, and analysis of valuable materials and enables applications such as digitizing of archives or assisting individuals with visual impairments [10].
This work focuses on the recognition of handwritten text at the line level, where the goal is to predict the character sequence contained in a single line of text. Recent work highlights this level of annotation as a practical and widely adopted choice in HTR, offering a balance between annotation effort and the ability to capture meaningful contextual information [17]. Compared to character- or word-level tasks, line-level recognition is more challenging and better reflects real document conditions [10].
HTR for Latin-script languages has reached near-human performance, especially for English [10]. In contrast, the same level of success has not been achieved for Arabic-script languages, which are used by more than 400 million people worldwide [28]. This gap is often attributed to the characteristics of Arabic script, which is cursive in nature, letters differing only by small diacritical marks, and shapes changing depending on their position within a word, and fewer, smaller, and less diverse datasets than for Latin-scripts [29]. To reduce this performance gap, research has mostly focused on architectural improvements while paying much less attention to data quality [6, 2].
@@
These issues do not arise randomly; they often originate from the semi-automated way in which handwritten text datasets are constructed, where automated methods are combined with human oversight. Because producing large datasets is time and labor intensive, this oversight is not always adequately managed, allowing samples with errors—such as incorrect or incomplete transcriptions, segmentation problems (e.g., truncated lines or multiple text lines within a single image), orientation mistakes, script-mismatched samples, or non-textual document elements (e.g., stamps or signatures)—to be included in the final dataset. Label errors are problematic as deep neural networks can memorize noisy labels during training , which harms generalization [33] and can distort model selection and benchmarking [22].
@@
Automated scoring alone can mistakenly flag valid but difficult samples as noisy. Prior work has shown that only about half of algorithmically flagged candidates correspond to genuine label errors [22]. Given the relatively small size of typical HTR datasets (e.g., IAM $\approx 10K$ lines [20], KHATT $\approx 6K$ [18]), human inspection of high-CER samples is practical. We therefore introduce CER-based Ranking with Human Verification (CER-HV), as a two-stage framework that combines CER-based ranking with human verification. To our knowledge, no prior work has combined noise detection with human verification in sequence-to-sequence settings such as HTR.
@@
    We establish a strong CRNN baseline for line-level Arabic-script HTR achieving state-of-the-art results on multiple datasets, including a reduction from 20.7% to 8.26% CER on Pashto (PHTI), and the first benchmarks for Ajami and Persian.
@@
    We provide the first publicly available cleaned evaluation splits and line-level benchmarks for Persian handwritten text (PHTD), including extracted line images and test partitions, enabling reproducibility and future benchmarking.
@@ 2.1 Arabic-script HTR
A similar direction is observed in Urdu HTR, where Transformer-based architectures have been used [26], often relying on additional printed or synthetic text data to supplement handwritten training sets due to their data-intensive nature [19, 13, 12].
@@ 2.2 Label Error Detection in Handwritten Text Recognition
Another family of methods exploits learning dynamics to identify noisy labels. O2U-Net ranks samples based on their average training loss over cycles of underfitting and overfitting, under the observation that mislabeled samples tend to maintain higher loss throughout training [14]. This idea is closely related to the memorization effect, where deep networks fit clean samples earlier than noisy ones. This early-learning phenomenon has been analyzed more broadly in the context of learning with noisy labels, where several works show that deep networks first fit clean labels before memorizing noisy ones, and propose training strategies that explicitly exploit or regularize this behavior to improve robustness [36]. While O2U-Net is formulated for classification, its reliance on per-sample loss trajectories makes it conceptually applicable to other settings. However, in HTR, the interpretation of per-sample CTC loss is complicated by alignment variability and structural errors, and cyclic overfitting strategies may not always be desirable. Prior work on early stopping shows that preventing late-stage memorization can be beneficial under label noise, suggesting that careful control of training dynamics is important when adapting such methods [32].
@@ 3.1 Problem Formulation
$P(y\mid x;\theta)=\sum_{\pi\in(\Sigma^{\prime})^{T},B(\pi)=y}\prod_{t=1}^{T}P(\pi_{t}\mid x;\theta).$ (2)
The model parameters $\theta$ are learned by minimizing the CTC loss over the training set
@@ 3.2 HTR Model: CRNN Architecture
[…]
@@
Figure 3: The CRNN architecture used as the base recognition model in Stage 1 of the CER-HV framework.
The CNN encoder extracts hierarchical visual features from the input text line image and converts the resulting two-dimensional feature maps into a one-dimensional sequence along the image width using column-wise max pooling. This preserves the horizontal reading order while providing a compact representation suitable for recurrent sequence modeling. The recurrent decoder then models contextual dependencies and produces per-frame character probabilities, which are decoded using CTC.
@@
Concretely, the CNN encoder begins with a $\times7$ convolution (32 channels, stride $\times2$), followed by ResNet blocks (2 with 64 channels, 3 with 128, and 2 with 256), interleaved with batch normalization and dropout. Progressive $\times2$ max-pooling between block groups downsamples the spatial resolution, producing feature maps of size (H/8) $\times$ (W/8) $\times$ 256. Column-wise max pooling then flattens this tensor into a sequence of (W/8) feature vectors (256 dimensions each).
@@ 3.3 CER-HV Framework: CER Based Ranking and HITL Verification
Although automated scoring highlights suspicious samples, some correctly labeled but visually difficult examples may also receive high CER. For this reason, a HITL verification step is used. Samples whose CER exceeds a threshold are reviewed:
@@
After human verification and cleaning, the cleaned dataset $D^{\prime}$ is used to retrain the model. The retrained recognizer is evaluated on the test set, and the same CER based scoring procedure can be used to identify remaining issues in the evaluation data. The process may be repeated if needed. Algorithm 1 summarizes the complete procedure. Beyond cleaning existing benchmarks, CER-HV can be applied as a general validation procedure to assess label quality in newly created or previously unverified line-level HTR datasets.
@@
 if sample has transcription, segmentation, orientation, or irrelevant content then
@@ 4 Experimental Setup
This section describes the datasets used, the training details, hyperparameters, input preprocessing, and experimental settings.
@@ 4.1 Datasets
This study employs six Arabic-script HTR datasets covering five languages segmented and labeled at the line level, covering both modern and historical handwriting. Together, these datasets provide broad coverage of writing styles, domains, and regional variations across Arabic, Persian, Urdu, Pashto, and West African Ajami scripts. Examples from the datasets are shown in Figure 4, and dataset statistics including train/validation/test splits are summarized in Table 1. For PHTD and Ajami without public train, validation, and test splits we applied an 80%, 10%, and 10% division. These datasets collectively enable assessment of the robustness of the proposed framework across diverse handwriting styles and linguistic varieties.
@@
KHATT (Arabic): The KHATT dataset [18] is a widely used benchmark for modern Arabic handwriting recognition. It contains paragraph images written by 1,000 writers from diverse demographic backgrounds. Each writer contributed both repeated and unique texts under controlled imaging conditions (black ink on white background). We used only the unique-text subset, yielding 6,742 segmented line images.
@@
NUST-UHWR (Urdu): The NUST-UHWR dataset [34] is the first published large-scale, unconstrained Urdu handwriting corpus spanning seven topical domains to ensure linguistic variety. It includes 10,608 text lines written by nearly 1,000 contributors.
@@ 4.2 Training Setup and Evaluation
For each dataset, image preprocessing followed a standardized resizing and padding procedure to reduce shape distortion and ensure uniform input dimensions. Specifically, all line images were resized to the rounded average height and width (in whole pixels) of their respective training set, as reported in Table 1. If an image exceeded these target dimensions, it was proportionally rescaled; otherwise, it was centered and padded with median pixel values. This approach preserved the aspect ratio while providing consistent spatial context across samples, which we observed stabilized training and improved recognition performance on longer text lines.
To further regularize training and enhance robustness to handwriting variability, we applied a consistent set of data augmentations across all datasets. These include affine transformations (small rotation and skew), grid distortion, elastic deformation, Gaussian noise addition, brightness and contrast adjustment, and morphological operations. Such augmentations have been shown to improve generalization in HTR tasks by encouraging invariance to writer style, line curvature, and ink intensity.
For each dataset, a separate CRNN model was trained for a maximum of 800 epochs with a batch size of 16 using the Adam optimizer and an initial learning rate of $5\times 10^{-4}$. A multi-step learning rate scheduler reduced the learning rate by a factor of 0.1 at 50% and 75% of the total epochs. Early stopping with patience $P=20$ epochs was applied based on validation CER to prevent overfitting. For each dataset, the model achieving the lowest CER on the validation set was selected for evaluation on the test partition.
@@ 5 Results
We begin by comparing our CRNN results with prior work across all datasets using the original, unaltered data to establish quantitative CER baselines. We then apply the CER-HV framework to analyze label noise across the training, validation, and test splits. Representative qualitative examples of common labeling issues are presented to illustrate the identified error categories. Finally, we quantify the impact of correcting these errors by evaluating how dataset cleaning influences recognition performance.
@@ 5.1 Quantitative Benchmarking of the CRNN-Based HTR System
Dataset (Language) Architecture Syn. CER (%)
KHATT (Arabic) [18] SFR [28] Yes 14.1
EfficientNetB4+BLSTM [5] Yes 13.34
FCN [37] No 8.77
CRNN (Ours) No 8.45 $\pm$ 0.14
Transformer (cross-attn.) [21] Yes 18.45
TrOCR [6] Yes 15.4
DAN [2] No 8.9
Muharaf (Arabic) [28] SFR [28] Yes 18.1
CRNN (Ours) No 10.11 $\pm$ 0.055
TrOCR [6] Yes* 11.7
PHTI (Pashto) [15] MD-LSTM [16] No 20.7
CRNN (Ours) No 8.26 $\pm$ 0.05
PHTD (Persian) [1] CRNN (Ours) No 11.3 $\pm$ 0.07
NUST-UHWR (Urdu) [34] Modified CRNN [34] No 7.35
Modified CRNN+LM [34] No 5.49
ET-Network (No Attention) [12] No 7.79
CRNN (Ours) No 6.66 $\pm$ 0.18
ViT+T5+GPT2 [13] No 17.0
CNN-Transformer [19] Yes 6.2
ET-Network (with Attention) [12] Yes 5.70
Conv-Transformer [26] No 6.4
Ajami (Fulfulde/Hausa) [38] Pre-trained models from OPEN ITI [38] No 64–84
CRNN (Ours) No 10.66 $\pm$ 0.04
Table 2: Comparison of our CRNN with established methods from the literature. Syn. indicates the use of synthetic or additional data. * The TrOCR result on Muharaf is reported as using additional 11K restricted samples in the original paper (marked as ‘Yes*‘ in the table). Results are reported as CER (mean $\pm$ standard deviation). Dashed lines separate non-Transformer and Transformer-based approaches.
@@ 5.2 Label Error Analysis
Using the CER-HV process described in Section 3.3, we analyzed label and content errors across all six datasets and across all dataset splits (training, validation, and test). The analysis reveals that the presence, type, and prevalence of labeling issues vary substantially depending on both the dataset and the split. In total, five datasets contained mislabeled or corrupted samples. PHTD did not exhibit any detectable labeling errors and is therefore excluded from the subsequent error analysis. For the remaining datasets, entries marked with “—” indicate that the corresponding error type was not observed for that dataset and split.
@@
Table 3: Precision of the CER-based noise detection over the top 50 highest-ranked samples in each split for datasets with sufficient numbers of flagged samples. Precision reflects the proportion of flagged samples that were confirmed as true label errors; the remaining cases correspond to valid but difficult samples.
@@
Dataset
Split
Transcription Error
Segmentation Error
Orientation Error
Script Mismatch
Irrelevant Content
@@
Table 4: Distribution of detected label errors by type across training, validation, and test splits.